OpenStudy (some.random.cool.kid):

A store had 100 t-shirts. Each month, 30% of the t-shirts were sold and 25 new t-shirts arrived in shipments. Which recursive function best represents the number of t-shirts in the store, given that f(0) = 100?

OpenStudy (some.random.cool.kid):

somebody help explain

OpenStudy (some.random.cool.kid):

anyone?

OpenStudy (anonymous):

so 100/30= 30 shirts sold, if 25 shirts come in you have 95 shirts

OpenStudy (some.random.cool.kid):

@SolomonZelman

OpenStudy (some.random.cool.kid):

can you help explain this

OpenStudy (some.random.cool.kid):

@SolomonZelman please help explain

OpenStudy (solomonzelman):

So, after the first months, we get, \(\large\color{black}{ f(1)= 100\times 0.7+25=70+25=95 }\) After the 2nd month we get, \(\large\color{black}{ f(2)= 95\times 0.7+25=66.5+25=91.5 }\) After the 3rd month we get, \(\large\color{black}{ f(2)= 91.5\times 0.7+25=64.05+25=89.05 }\)

OpenStudy (solomonzelman):

the last one is f(3), not f(2)

OpenStudy (some.random.cool.kid):

ok

OpenStudy (solomonzelman):

I don't see any recursive formula that can be written for any f(x) starting from f(0), but I can relate each a(n-1) and a(n). \(\large\color{black}{ a_n=a_{n-1}\times0.7~~~+25 }\)

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

so how exactly do i solve this though

OpenStudy (solomonzelman):

What do you mean "solve" ? what exactly do you have to solve for?

OpenStudy (some.random.cool.kid):

Which recursive function best represents the number of t-shirts in the store, given that f(0) = 100?

OpenStudy (some.random.cool.kid):

thats what im solving

OpenStudy (solomonzelman):

I gave you \(\large\color{blue}{ a_n=(a_{n-1}) \times 0.7~~+25 }\) from one term to the next one. There is really no rule that can represent the way you go from \(\large\color{blue}{ a_1 }\) till any \(\large\color{blue}{ a_n }\) as you would normally get by a sequence.

OpenStudy (some.random.cool.kid):

oh ok thanks

OpenStudy (some.random.cool.kid):

bye enjoy your day.

OpenStudy (solomonzelman):

you too, yw:)